Three equal charges 1.8*10^(-6) C each are located at the corners of an equilateral triangle ABC of each side 10 cm. Calculate the electric potential due to these charges at the midpoint of AB.
1
Expert's answer
2019-05-20T14:18:18-0400

The potential of a charge e at a distance d is given by


"\phi = \frac{ke}{d}"


with k = 8.988 109 V m / C.

If the length of a side of the triangle is l, the corresponding distances are 0.5l, 0.5l for the charges in points A, B and

"\frac{\sqrt{3}}{2} l"

for the charge in point C due to known expression for the altitude.

Thus


"\phi = \phi_A + \phi_B + \phi_C = \frac{2ke}{\,\,l\,\,} + \frac{2ke}{l} + \frac{2ke}{\sqrt{3}l}"

"\phi = \frac{ke}{l} (4+\frac{2}{\sqrt{3}}) \approx 8.3\cdot 10^{5} V"